Calculate the hydronium concentration of a certain ionization reaction where the hydroxide concentration is determined to be 2.7 x 10-5 M. Identify if the solution is acidic, basic, or neutral.


1
Expert's answer
2020-11-26T05:09:57-0500

pOH= -log(OH-)

=-log(2.7×10^-5)

=4.569

pH= 14-pOH

pH= 14-4.569

=9.431

[H30+]= antilog(-pH)

=antilog(-9.431)

=3.707×10^-10 M

The solution is basic since pH is 9.431 and the concentration of hydroxide ions is more than that of hydronium.